OpenStudy (anonymous):

One apartment is directly above a second apartment. The resident living downstairs calls his neighbor living above him and states, " If one of you is willing to come downstairs, we'll have the same number of people in both apartments." The upstairs resident responds, " We're all too tired to move. Why don't one of you come up here? Then we'll have twice as may people up here as you've got down there." How many people are in each apartment? *PLEASE SHOW YOUR WORK*

OpenStudy (anonymous):

u're kidding right??

OpenStudy (anonymous):

no, you're kidding. I am dead serious!

OpenStudy (anonymous):

sorry i'm not a geek..:P

OpenStudy (anonymous):

right...so i got answers which match up.  so your first eq will be:---> x - 1 = y + 1 [i] and your second one will be:----> x + 1 = 2(y -1) [ii] * i just made the two eq's from the statements above so...from the first equation you isolate a variable, lets say x so to make x the subject, you get the equation x = y + 2 and substitute it into the second one to give you:---> y+3 = 2y -2 and you get y = 5 replace y in the first equation and you'll get x = 7 so...the top apartement has 7 people and the bottom one has 5. Tell me if you want more explanation.